OpenStudy (anonymous):

PLEASE HELP ANSWER!! Given f(x) = –3x – 7, what is f(–4)? A. 40 B. 5 C. –1 D. –14

OpenStudy (hesan):

Simply Replace x by -4... You will get the answer..

OpenStudy (anonymous):

no

OpenStudy (hesan):

And that is B. 5

OpenStudy (anonymous):

thank you!

OpenStudy (lieutenantgeneral):

f(x) = -4 -3 * -4 = 12 12 - 7 = 5 B.

OpenStudy (anonymous):

thanks!